Calculate the acceleration of the electron which electronic charge is given as (1.60×10^-19c) and mass of electron as(9.1×10^-3kg)


1
Expert's answer
2021-09-02T16:46:11-0400

Let's first find the electric force that acts on the electron:


"F=qE."

Unfortunately, from the definition of the question we don't know the magnitude of the electric field, "E". Let's assume that "E=200\ \dfrac{N}{C}." Than, we can find the electric force:


"F=qE=1.6\cdot10^{-19}\ C\cdot200\ \dfrac{N}{C}=3.2\cdot10^{-17}\ N."

Finally, we can find the acceleration of the electron from the Newton's Second Law of Motion:


"a=\dfrac{F}{m}=\dfrac{3.2\cdot10^{-17}\ N}{9.1\cdot10^{-31}\ kg}=3.5\cdot10^{13}\ \dfrac{m}{s^2}."